OpenStudy (anonymous):

obtain in cartesian form all values of logz z = -3-2i

OpenStudy (anonymous):

\[\log(z)=\ln(|z|)+i\text{arg}(z)\] \[|-3-2i|=\sqrt{3^2+2^2}=\sqrt{13}\] \[\arg(z)=\tan^{-1}(\frac{2}{3})+\pi\]

OpenStudy (anonymous):

thank u